Invitation to Reception Party (Takeshi Hirai)

Invitation to Reception Party on Thursday, the 21st February, Evening 18:00 – 20:30 prepared by Mr. and Mrs. HIRAI,

Dear Collleagues, Participants of the workshop held at RIMS, 18th – 22nd, February

We, Mr. and Mrs. HIRAI, will prepare a reception party for all the participants of this workshop, in the evening of Thursday, the 21st February, in the campus of Faculty of Integrated Human Studies.

We cordially invite you to this informal party, and we will be very happy if you could join us.

No formal speaches are assumed, and you can just enjoy freely foods, drinks and conversations.

The details are as follows.

Date and Time: Thursday, the 21st February, Evening 18:00 〜 20:30

Place: 4th floor, Seminar Room #1401(Room Tel: 753-7895) in the building No.1 at the left side of the main gate Caution: The doors of the entrance of this building will be closed at about 7 o’clock. After this time you cannot open the door from the outside (it can be opened from the inside).

In case you have some problems, please call the phone number indicated above.

13/02/2002

Takeshi and Etsuko Hirai